OpenStudy (anonymous):

Please Help! Are the functions linear or nonlinear? 1. 72=x^3+y 2. y+1=5(x-9) 3. 7y+2x=12 4. 4y=24

OpenStudy (anonymous):

solve each of the equations for y, so that you get the equation for a line. 1) y = -x^3 + 72 2) y = 5x - 46 3) y = (-2/7)x + (12/7) 4) y = 6

OpenStudy (anonymous):

Okay! thanks